rms value of i= 10 sinwt + 20 cos (wt + 30°)

1
Expert's answer
2021-12-02T20:19:58-0500

The rms value by definition

"i_{\rm rms}=\sqrt{\frac{1}{T}\int_{0}^{T}i^2(t)dt}"

"\int_{0}^{T}(10\sin (2\pi t/T) + 20\cos (2\pi t/T + 30°))^2dt"

"=150T"

Finally

"i_{\rm rms}=\sqrt{150}=12.2"